OpenStudy (anonymous):

Let f(x) = x^2 – 16. Find f–1(x)

OpenStudy (abb0t):

solve for x in terms of y.

OpenStudy (anonymous):

yes

OpenStudy (abb0t):

No. I mean that's how you find the inverse. lol

OpenStudy (anonymous):

i know, but i'm not sure how to exactly do it

OpenStudy (abb0t):

\[x = y^2-16\] You can add 16 to both sides to get: \[x+16=y^2\] then take the square root.

OpenStudy (anonymous):